a house wife average expenditure during a week was Rs 35 for the first four days and Rs 38 for the next two days .if her average expenditure for the day of the week was Rs 36,how much did she spend on the seventh day?
Answer by rfer(16322)

(4*35+2*38+x)/7=36
140+76+x=36*7
x=252-216
x=36